Some eye surgeries temporarily insert a tiny 0.125 mL gas bubble, pressurized to standard 101.3 kPa, to keep the retina in place during healing. However, patients are warned not to scuba dive, since pressure can reach 4.5 times that of standard pressure at typical diving depths. Calculate the volume, in mL, that the gas bubble will obtain at typical diving depths.


1
Expert's answer
2021-02-05T04:07:06-0500

Initial Volume(V1) = 0.125 mL

Intial Pressure (P1) = 101.3 kPa

Final pressure (P2) = 4.5 × 101.3 = 455.85 KPa

Using Boyle's law,

P1V1 = P2V2

101.3 × 0.125 = 455.85 × V2

V2 = 0.0312 mL

Volume = [0.0312] mL